### v1

# OpenFake: An Open Dataset and Platform Toward Large-Scale Deepfake Detection

Abstract
Deepfakes, synthetic media created using advanced AI techniques, have intensified the spread of misinformation, particularly in politically sensitive contexts. Existing deepfake detection datasets are often limited, relying on outdated generation methods, low realism, or single-face imagery, restricting the effectiveness for general synthetic image detection. By analyzing social media posts, we identify multiple modalities through which deepfakes propagate misinformation. Furthermore, our human perception study demonstrates that recently developed proprietary models produce synthetic images increasingly indistinguishable from real ones, complicating accurate identification by the general public. Consequently, we present a comprehensive, politically-focused dataset specifically crafted for benchmarking detection against modern generative models. This dataset contains three million real images paired with descriptive captions, which are used for generating 963k corresponding high-quality synthetic images from a mix of proprietary and open-source models. Recognizing the continual evolution of generative techniques, we introduce an innovative crowdsourced adversarial platform, where participants are incentivized to generate and submit challenging synthetic images. This ongoing community-driven initiative ensures that deepfake detection methods remain robust and adaptive, proactively safeguarding public discourse from sophisticated misinformation threats.
1 Introduction
Figure 1: We begin by scraping politically relevant images from social media (e.g., X, Reddit, Bluesky), filtered by election-related hashtags. Manual investigation of these social media images helps us to design a prompt for filtering politically relevant images. A vision-language model (e.g., Qwen2.5-VL) extracts thematic captions or prompts from real images from LAION. These prompts serve dual purposes: (1) forming a large bank of real image–prompt pairs, and (2) seeding generation across a range of synthetic image models (e.g., SDv3.5, Flux, Ideogram, GPT Image 1).
Deepfakes, realistic synthetic media generated by AI, have emerged as a serious threat to the information ecosystem (Canadian Security Intelligence Service, 2023; Bengio et al., 2025c). By enabling anyone to fabricate audio-visual content of real people, deepfakes can spread false information at an unprecedented scale, eroding trust across various platforms, from social media and online content to traditional media outlets. High-profile cases (e.g., forged speeches or imagery of public figures) and the prevalence of non-consensual intimate imagery underscore the potential for harm to political stability, reputation, and public safety Marchal et al. (2024). Scholars have warned of an “infopocalypse” where constant exposure to fake media breeds cynicism or paranoia Schick (2020). Detecting deepfakes reliably is therefore critical to mitigate the spread of misinformation and disinformation [^1] We adopt the term misinformation throughout this paper to refer broadly to harmful or misleading content. Technically, misinformation denotes false information shared without intent to deceive, while disinformation refers to deliberately deceptive content. Our usage includes both, given the difficulty of inferring intent in large-scale media analysis., and to restore trust in digital media. The rapid advancement of AI-generated image technologies has reached a point where distinguishing between real and synthetic images has become increasingly challenging for humans. Studies have shown that humans underperform in identifying AI-generated images, highlighting the sophistication of these generative models (Diel et al., 2024).
The political sphere is particularly vulnerable to the risks posed by deepfakes, which can be weaponized to manipulate public opinion and undermine democratic processes (Bengio et al., 2025c, b; Karen Hao, 2019). Synthetic media have already been exploited for scams, blackmail, and targeted reputation sabotage, while the fabrication of fake historical artifacts, manipulated medical images, and staged events introduces new avenues for the spread of misinformation and societal harm (Ferrara, 2024; Bengio et al., 2025c). By flooding social and traditional media with convincing falsehoods, deepfakes erode public trust in news and create confusion about what is real, particularly during sensitive periods like elections (IVADO and CEIMIA, 2025). Such disruptions threaten not only individual reputations and public safety, but also the legitimacy of democratic institutions and processes, with scholars warning that advances in generative AI could empower malicious actors to influence political outcomes and destabilize societies (Bengio et al., 2025a; Hameleers et al., 2024; Vaccari and Chadwick, 2020).
Based on this political motivation, we built our dataset by starting with a large-scale collection of politically relevant social media posts, including tweets, images, and videos scraped during the 2025 Canadian federal election. While this initial focus helped identify image types likely to appear in high-impact or misleading contexts, the final dataset spans a broad range of content beyond political imagery, including everyday scenes, people, and objects. These posts serve as a foundation for analyzing real-world misinformation risks and curating image content with societal relevance. However, relying directly on social media data introduces challenges such as label contamination and unreliable ground truth, which we sought to avoid in contrast to prior work (Chen and Zou, 2023). To assess the deception risk of current generative models, we also conducted a user study measuring people’s ability to distinguish real from synthetic political images, revealing that outputs from some proprietary models may approach chance-level accuracy.
Despite significant progress in deepfake detection research, current datasets suffer from major limitations that restrict their effectiveness in modern, real-world scenarios. As shown in Table˜1, most established benchmarks rely on outdated generation methods; GAN-based face-swapping tools such as DeepFaceLab (Liu et al., 2023) and Face2Face (Thies et al., 2016). These datasets, while valuable for early detection efforts, are increasingly unrepresentative of the latest synthetic media, particularly high-fidelity diffusion and transformer-based models. Moreover, they overwhelmingly focus on single-face portraits, providing little to no real-world grounded images and neglecting the broader spectrum of image-based misinformation that floods political and social media discourse: crowd scenes, protests, disaster images, manipulated signage, or synthetic screenshots.
To address these gaps, we introduce OpenFake [^2] https://huggingface.co/datasets/ComplexDataLab/OpenFake, a politically grounded dataset designed for general deepfake detection. OpenFake focuses on synthetic media that poses genuine misinformation risks. It pairs large-scale real image corpora with state-of-the-art synthetic images. A high-quality dataset like this helps benchmark existing detectors and supports the development of stronger ones. Generative and detection models evolve in a feedback loop: as new generative methods are released, new detectors must be built, which can then inform on directions to improve the generators even further. As generation methods advance rapidly, static datasets risk becoming outdated. To ensure our dataset remains relevant over time, we introduce OpenFake Arena, a crowdsourced adversarial platform where users are invited to generate synthetic images that fool a live deepfake detection model. Successful submissions are validated using a CLIP-based prompt similarity gate, scored against a continually updated classifier, and, if they pass, contributed back into the dataset, forming a self-improving benchmark. This enables continual learning and ensures the benchmark remains both current and increasingly challenging.
In summary, our contributions are:
  • A large, politically relevant dataset of 3 million real images paired with extracted prompts, curated for misinformation risk and motivated by social media content.
  • A diverse, high-quality synthetic image set spanning 963k samples, generated from both open-source and proprietary models.
  • A human perception study revealing that some modern proprietary generators can reduce human classification accuracy to nearly random chance.
  • Baseline analysis shows detectors trained on OpenFake outperform those from older datasets against high-quality deepfakes, showing its value for modern detection.
  • A scalable crowdsourcing framework (OpenFake Arena) for adversarial image generation, enabling continual community-driven benchmarking.
Together, OpenFake and OpenFake Arena form a robust and adaptive foundation for studying deepfakes in politically sensitive contexts, providing researchers and practitioners with the publicly available tools necessary to characterize emerging synthetic threats.
2 Related work
Synthetic image datasets.
Despite the proliferation of generative models, existing deepfake datasets remain limited in realism, diversity, and accessibility. Early benchmarks such as FaceForensics++ (Rössler et al., 2019), Celeb-DF (Li et al., 2020), and DFDC (Dolhansky et al., 2020) rely on GAN-based face-swapping techniques and focus almost exclusively on single-person portrait videos. Even newer datasets such as ForgeryNet (He et al., 2021), OpenForensics (Le et al., 2021), and FFIW (Zhou et al., 2021) continue to emphasize face-centric detection, with limited variation in image content or generation method (Cheng et al., 2024; Chen et al., 2024b; Yan et al., 2024). More recent image datasets have started to incorporate diffusion-based generators (e.g., Stable Diffusion, DALL·E 2, Midjourney), as seen in Fake2M (Lu et al., 2023), DiffusionForensics (Wang et al., 2023), and GenImage (Zhu et al., 2023). However, these datasets still fall short in several ways. First, most rely on open-source models like SDv1.5 or SDv2.1 (Rombach et al., 2022), which, while important, do not match the visual fidelity of cutting-edge proprietary models such as Imagen 3 (Baldridge et al., 2024) or GPT Image 1 (OpenAI, 2025). As a result, they fail to represent the modern threat landscape posed by the most deceptive fakes. Second, many datasets lack real-world grounding. Image prompts are frequently abstract, artistic, or class-based (e.g., GenImage uses classes from ImageNet-1k (Deng et al., 2009)), failing to capture the multimodal misinformation strategies actually deployed online. Third, these datasets are static and infrequently updated, meaning they quickly become outdated as generation tools evolve. Fourth, prompts used for image generation are often withheld, making it difficult for others to reproduce, regenerate, or expand these datasets with future models. In contrast, we release a large bank of extracted prompts along with the images, which enables researchers to extend the dataset using alternative text-to-image models while preserving topical consistency. Finally, accessibility remains a persistent issue. Many datasets require downloading large zip archives via Google Drive or web links, making them difficult to integrate into new pipelines. In contrast, OpenFake is fully hosted on the HuggingFace Hub in streaming-friendly Parquet format, enabling scalable access and evaluation, which should help the community develop new detection tools. Table˜1 highlights these differences in model coverage, dataset scope, prompt extensibility, and access modality.
Dataset Year # Fakes # Reals Extendable Content Scope Realism Access # Methods Most recent model
FaceForensics++ (Rössler et al., 2019) 2019 5K 1K  Narrow Low  Hard 4 Face2Face (2016)
Celeb-DF (Li et al., 2020) 2020 5K+ 590  Narrow Low  Hard 1 DeepFaceLab (2020)
DFDC (Dolhansky et al., 2020) 2020 100K+ 20K+  Narrow Low  Hard 8 DeepFaceLab (2020)
ForgeryNet (He et al., 2021) 2021 1.5M 1.5M  Narrow Low  Hard 15 DeepFaceLab (2020)
FFIW (Zhou et al., 2021) 2021 10K 10K  Narrow Low  Hard 3 DeepFaceLab (2020)
OpenForensics (Le et al., 2021) 2021 100K 100K  Narrow Low  Hard 3 GAN (2020)
DeepFakeFace (Song et al., 2023) 2023 90K 30K  Narrow Medium  Hard 3 SD v1.5 (2022)
Fake2M (Lu et al., 2023) 2023 2M 0  Moderate Medium Easy 3 SD v1.5 (2022)
DiffusionForensics (Wang et al., 2023) 2023 570K 140K  Moderate Medium  Hard 8 iDDPM (2021)
DMDetection (Corvi et al., 2023) 2023 200K 200K  Moderate Medium  Hard 3 DALL·E 2 (2022)
GenImage (Zhu et al., 2023) 2023 1.3M 1.33M  Moderate Good  Hard 5 Midjourney 5 (2023)
TWIGMA (Chen and Zou, 2023) 2023 800K 0  Rich Medium  Unavailable
DiffusionDeepfake (Bhattacharyya et al., 2024) 2024 100K 94K  Narrow Good  Hard 2 Midjourney (2024)
DF40 (Yan et al., 2024) 2024 1M+ 1.5K  Narrow Medium  Hard 40 PixArt-$\alpha$ (2024)
DiffusionFace (Chen et al., 2024b) 2024 600K 30K  Narrow Good  Hard 11 SD v2.1 (2022)
DiFF (Cheng et al., 2024) 2024 500K 23K  Narrow Good  Hard 13 Midjourney 5 (2023)
Semi-Truths (Pal et al., 2024) 2024 1.34M 26K  Moderate Good Easy 8 Stable Diffusion XL (2023)
OpenFake (Ours) 2025 963K 3M  Rich High Easy 18 Imagen 4.0 (2025)
Table 1: Overview of public deepfake datasets. “# Fakes” and “# Reals” count individual media items (images or videos; units omitted for brevity). Content Scope:  Narrow (face-focused);  Moderate (diverse but limited);  Rich (broad, internet-like variety). Access:  Unavailable;  Hard (public but cumbersome);  Easy (clean, ready-to-use hosting). Realism: qualitative fidelity of synthetic content — Low, Medium, Good, High. Extendable: availability of prompts/metadata enabling seamless dataset expansion. OpenFake uniquely combines rich scope, high realism, large real sample count, easy access, and extensibility.
Deepfake detection methods.
Early detection approaches relied on convolutional neural networks trained on known forgery artifacts, such as blending boundaries or low-level inconsistencies (Afchar et al., 2018; Rössler et al., 2019; V and Joy, 2023). While effective in-domain, these models struggle to generalize across generation techniques (Ojha et al., 2023). As diffusion and transformer-based models reduce such artifacts, recent work has focused on semantic-level signals and frequency-domain cues (Durall et al., 2020; Liu et al., 2021a; Frank et al., 2020; Qian et al., 2020). CLIP-based detection (Cozzolino et al., 2024; Khan and Dang-Nguyen, 2024) has emerged as a promising direction, leveraging large-scale vision-language pretraining to improve robustness. Other advances include domain-adaptive feature learning (Shao et al., 2025; Jia et al., 2024), zero-shot detectors (Lin et al., 2024), and hybrid approaches that blend local artifact patterns with global semantic reasoning (LI et al., 2024; Zhou et al., 2024; Ma et al., 2025). Despite progress, the rapid pace of generative model development continues to outstrip detection capabilities, motivating adaptive benchmarks like OpenFake Arena to assess robustness in a dynamic, adversarial setting.
3 Empirical Motivation: Real-World Misinformation and Human Limits
3.1 Social media content
Social media platforms, notably X (formerly Twitter), have become critical channels for political discourse, and consequently, for amplifying deepfake disinformation. To better understand the real-world impact of deepfakes, we conducted an analysis of images shared on X.
(a) Obvious fake
(b) Text embedded
(c) Realistic faces
(d) Realistic sign
(e) News event
Figure 2: Examples of deepfake images collected from X depicting various types of fabricated scenarios involving Canadian political figures and events.
As part of a subsequent study, we have collected social media posts—including images and videos—using a cloud-based, real-time multimedia data collection pipeline deployed on AWS. This pipeline continuously monitors and filters public content across three platforms: X, Reddit, and Bluesky. The data capture spans the period leading up to and including the 2025 Canadian federal election, enabling analysis of online discourse during this critical time.
Systematic manual verification of 650 randomly sampled images collected during a 24-hour period identified 51 distinct deepfake instances (7.8%). Verification methodology involved source authentication and reverse image search techniques. This detection rate should be considered conservative, as the actual prevalence of deepfakes within the dataset is likely higher due to limitations in manual detection methods. Notably, many of these deepfakes prominently featured political figures. For instance, fabricated scenarios involving leading candidates significantly distorted their public perception. Even when certain deepfakes were clearly artificial (Figure˜2 a), featuring visual inconsistencies or unrealistic scenarios, they effectively reinforced existing biases or narratives, resonating more profoundly with audiences compared to textual misinformation alone (Ecker et al., 2022; Hameleers et al., 2020; Vaccari and Chadwick, 2020). When the deepfake is convincing, the potential for harm increases substantially. Viewers are more likely to accept the content as genuine, especially when the image aligns with their preexisting beliefs or emotional biases. For example, as shown in Figure˜2c, a photorealistic yet entirely fabricated image of a political figure in an implied context, despite lacking any overt manipulation cues, can mislead viewers through association alone.
Beyond fabricated portraits of politicians, misinformation via images also extended to signs, banners, and manipulated visuals of significant news events such as protests or catastrophes as seen in Figure˜2 d) and e). These images often appeared alongside carefully crafted textual content designed to reinforce misleading narratives as exemplified in Figure˜2 b). Advanced Vision-Language Models (VLMs) enable the extraction of any textual content present in images, so our dataset does not need to explicitly target images with visible text. Any such content can still be leveraged for multimodal misinformation detection. Prior work has shown that combining VLM-extracted text with Large Language Model (LLM)-based fact-checkers can enhance automated verification systems (Tahmasebi et al., 2024; Raza et al., 2025; Li et al., 2024).
Altogether, our analysis highlights that deepfakes, while often focused on portraits, extend far beyond facial imagery. They frequently depict fabricated news events, disaster scenes, protest visuals, and manipulated political symbols—forms of visual misinformation that can be equally, if not more, influential. These findings directly inform the design of our dataset: to support generalizable detection, a deepfake benchmark must move beyond faces to capture the full breadth of misleading visual content. This motivates the use of a targeted filtering pipeline and underscores the need for robust, scalable detection systems equipped to handle the multimodal nature of modern misinformation.
3.2 Human Perception Study
Real GPT Image 1 Imagen 3 Grok SDv2 SDv3.5 Flux.1-dev
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Examples of deepfakes from each model used in the survey with their respective real image.
To assess the difficulty of detecting deepfakes generated by different models, we conducted a simple human study. A total of 100 participants completed the survey, each viewing a randomized set of 24 images. The set consisted of 12 real photographs and 12 synthetic images—2 from each of the six generation models: GPT Image 1 (OpenAI) (OpenAI, 2025), Imagen 3 (Google) (Baldridge et al., 2024), Grok 2 (xAI), Flux.1.0-dev (Black Forest Labs) (Labs, 2024), Stable Diffusion 2.1 Rombach et al. (2022), and Stable Diffusion 3.5. All synthetic images were generated from the same prompts as their real counterparts, using the text automatically extracted by our pipeline described in Section˜4. We did not hand-pick or optimize the generations, aside from removing names and content restrictions when necessary. Each prompt was only shown once to a given participant. This ensured that responses reflected a fair and diverse exposure across the dataset. In total, the survey contained 168 unique images. The survey description, code, and images are available at https://github.com/vicliv/deepfake-survey.
Model
Release
Access type
Humans
CLIP-D-10k+
Corvi2023
Fusion
(CLIP+Corvi)
SwinV2
SwinV2
(Semi-Truths)
SwinV2
Real 0.718 0.479 0.000 0.062 1.000 0.500 1.000
Imagen 3 2024 Proprietary 0.490 0.458 0.708 0.667 0.625 1.000 1.000
GPT Image 1 2025 Proprietary 0.684 0.458 0.500 0.458 0.750 1.000 1.000
Flux.1.0-dev 2024 Open-source 0.689 0.562 0.792 0.812 0.917 1.000 1.000
SDv3.5 2024 Open-source 0.709 0.521 1.000 0.938 0.750 1.000 1.000
Grok 2 2024 Proprietary 0.811 0.771 1.000 0.938 0.208 1.000 0.875
SDv2.1 2022 Open-source 0.879 0.646 1.000 0.938 0.917 1.000 1.000
0.714 0.524 0.417 0.427 0.804 0.750 0.990
Table 2: Survey results showing human true positive rate (TPR) for synthetic images and true negative rate (TNR) for real images. A score near 0.50 indicates chance-level performance. SwinV2 trained on our data achieves high accuracy on detecting fakes from the models it was trained on, but underperforms on unseen or low-quality deepfakes. Grok 2 is kept out of the training data to verify SwinV2’s out-of-distribution performance (only for the survey results).
The results in Table˜2 reveal several key insights into the current landscape of synthetic image realism and human perception. Most notably, Imagen 3 from Google achieved the lowest human accuracy (48.5%), performing no better than random guessing. GPT Image 1, OpenAI’s recent image generation model, also ranked among the most deceptive, with nearly one-third of its images going undetected. By contrast, Stable Diffusion 2.1 (SDv2.1), which is widely used in current benchmarks, had the highest human detection rate (87.9%) and was easily flagged by all detectors. This suggests that while open-source models are still relatively easy for both humans and machines to spot, advanced proprietary models exhibit an exceptionally high degree of realism and consistency, underscoring the importance of including them in benchmarks when evaluating the limits of human and machine perception. This demonstrates that proprietary models are setting a new bar in generation quality, one that current deepfake detection benchmarks, largely trained on older or open-source models, fail to represent.
However, this perceptual difficulty is not limited to humans. Baseline deepfake detectors also failed to consistently identify these advanced fakes. The CLIP-D-10k+ method (Cozzolino et al., 2024), which fine-tunes a linear classifier on CLIP embeddings, performed close to random on several models (e.g., 45.8% on Imagen 3, and GPT Image 1), and failed to distinguish real images entirely (47.9% TNR). Similarly, the Corvi2023 method (Corvi et al., 2023), which uses curated handcrafted features for detection, fared better on some open-source models (e.g., 100% on SDv3.5), but completely failed on real images (0.0% TNR) and newer proprietary content like GPT Image 1 (50.0%). In contrast, our SwinV2 baseline, trained directly on a curated mix of real and synthetic images from models included in our dataset, achieved 99.0% overall accuracy and perfect performance on its in-distribution models while having a reasonable score of 87.5% on Grok 2 (kept out of the training for this test). While this result demonstrates that deep networks can learn to detect even the most realistic fakes with sufficient supervision, it also highlights the brittleness of existing methods when faced with novel or unseen generative sources. The two additional SwinV2 baselines confirm this pattern: models trained on smaller or differently distributed data (GenImage and Semi-Truths) showed weaker generalization to modern deepfakes, with accuracy dropping to 80.4% and 75.0%, respectively. These results reinforce the importance of training on a large, diverse dataset like OpenFake to ensure robustness against emerging generation techniques.
In summary, both humans and the older detection systems we tested struggle to identify deepfakes generated by modern diffusion and transformer-based models. These findings call into question the utility of outdated benchmarks and reinforce the need for detection datasets that reflect the state of the art. Our study motivates a shift toward more diverse, high-fidelity, and license-aware synthetic corpora, paired with regular human benchmarking, to ensure that future detectors remain robust in the face of rapid generative progress.
4 Dataset Overview & Collection
Figure 4: t-SNE visualization of CLIP vision embeddings for 3,000 test images, including both real and synthetic images from a few generative models. Each point corresponds to an individual image, and colours indicate the generative model (or “real” for authentic images).
OpenFake combines a large repository of real images with a diverse collection of high-quality synthetic counterparts generated by multiple state-of-the-art models. In Appendix A, Table˜5 presents in details the key statistics of the dataset. Figure˜4 offers a qualitative view of the underlying distribution of some of the images. The substantial overlap between real and synthetic samples in the CLIP feature space highlights their semantic alignment, suggesting that synthetic images effectively mimic the distribution of real-world content. This shows that the prompt generation pipeline is working as intended. An overview of the data collection and generation process is provided in Figure˜1.
Real images.
We extract metadata from the LAION-400M dataset (Schuhmann et al., 2021), which we selected due to its broad representation of internet-sourced images—the same domain where visual misinformation typically circulates. Additionally, this dataset was likely included in the training data of the text-to-image models used to generate the synthetic images, which should theoretically make it more difficult for detectors to distinguish between real and fake images. More importantly, these images preserve real-world compression artifacts, which are crucial for training detectors that operate in the wild. While LAION may contain some synthetic images, we expect this contamination to be minimal, as the dataset primarily consists of content from 2014–2021, before the public release of diffusion models in 2022. After scraping LAION, we filter image–caption pairs using a vision–language model (Qwen2.5-VL (Bai et al., 2025)). As described by the prompt used to query the model in Appendix˜B, an image is retained if it is identified as depicting either (i) real human faces or (ii) politically salient or newsworthy events. For every retained image, we generate a more detailed caption to use as prompt input for text-to-image models. These $3\,\mathrm{M}$ prompts are also publicly released and form the basis of the prompts shown to users of our crowdsourcing platform OpenFake Arena.
We filtered real images using Qwen2.5-VL, selected for its trade-off between speed and quality. To prevent detection shortcuts, we excluded LAION images smaller than $\times512$ pixels from the released train/test sets, as lower resolutions introduced compression artifacts that made detection artificially easier. Full prompts used for filtering and captioning are provided in Appendix˜B, and additional details on generation and compute resources are in Appendix˜D.
Synthetic images.
We generated images from a diverse set of state-of-the-art generators using samples from our prompt bank: Stable Diffusion 1.5/2.1/XL/3.5 (Rombach et al., 2022), Flux 1.0-dev/1.1-Pro/Schnell (Labs, 2024), Midjourney v6/v7 (Midjourney, 2024, 2025), DALL·E 3 (OpenAI, 2023), Imagen 3/4 (Baldridge et al., 2024; Google Cloud, 2025), GPT Image 1 (OpenAI, 2025), Ideogram 3.0 (Ideogram AI, 2025), Grok-2 (xAI, 2024), HiDream-I1 (HiDream-ai, 2025b, a), Recraft v3 (Recraft, 2024), Chroma (Lodestones, 2025), and 10 community variants (Finetuned or LoRA) of Stable Diffusion 1.5/XL and Flux-dev. All images are produced at $\sim$1 MP resolution with varied aspect ratios (9:16, 16:9, 1:1, 2:3, 3:4, etc.), mirroring common social-media formats. Because several proprietary sources impose “non-compete” clauses, those subsets are released under a non-commercial license.
Splits and accessibility.
We construct the train/test split by sampling 1,000 images per generative model (with the exception of out-of-distribution models, which contribute between 200 and 600 images) along with the corresponding number of real images, yielding a test set of roughly 60,000 images (about 3% of the dataset assuming balanced classes). The remaining images are allocated to the training set. To ensure balance, each model is equally represented in the test split, and real images are matched accordingly. The rest of the real images and prompts are provided in a CSV file, with the real images accessible through their URLs. As OpenFake Arena expands and more synthetic images are collected, additional real images will be incorporated to preserve parity between real and synthetic domains in the train/test splits. All assets are hosted on the HuggingFace Hub at ComplexDataLab/OpenFake. All images have their associated prompts and model name as metadata, which can be used for model attribution.
Limitations.
Our dataset reflects limitations from its sources. Real images from LAION are skewed toward Western, pre-pandemic content, and generative models carry aesthetic and cultural biases. We mitigate risks via filtering, Arena submission review, and safeguards against misuse, but adversarial content and prompt noise remain challenges. More details are in Appendix˜C.
5 Crowdsourced Adversarial Platform
Generative and detection models co-evolve: advances in generation demand stronger detectors, which in turn promote new generation models. To keep benchmarks relevant amid rapid progress, we introduce OpenFake Arena: a crowdsourced platform where users generate synthetic images to fool a live detector. Successful examples are added to the benchmark, enabling sustained evaluation.
OpenFake Arena [^3] https://huggingface.co/spaces/CDL-AMLRT/OpenFakeArena is designed as a web-based interactive game to encourage wide participation. Each round begins with a prompt sampled from our bank of over 3 million. Users respond by generating a synthetic image using any generative model or editing tool that aligns with the prompt. A CLIP-based similarity gate verifies prompt-image alignment. If the image passes this check, it is evaluated by a detector trained on the OpenFake dataset. If the detector misclassifies the synthetic image as real, the user earns a point and the image is added to the benchmark.
The Arena features a real-time leaderboard to gamify the experience and incentivize participation. The detector is periodically retrained with newly collected data, enabling continual improvement. Submitted images are periodically reviewed. This human-in-the-loop setup transforms model drift from a challenge into a feature, allowing the benchmark to evolve organically alongside the state of generative models. Implementation details and screenshots of the arena are in Appendix˜E.
6 Baseline Detector Benchmarks
We evaluate a selection of deepfake detectors on the OpenFake dataset, with the goal of assessing how well existing models generalize to modern synthetic media, especially high-quality images from diffusion and transformer-based models.
Benchmark models.
We benchmark a broad set of representative detection paradigms, spanning supervised, semi-supervised, and zero-shot approaches. Our primary supervised detector is SwinV2-Small (Liu et al., 2021b), a vision transformer pretrained on ImageNet-1k (Deng et al., 2009) and finetuned on our dataset. We further evaluate two SwinV2 variants trained on external datasets (GenImage and Semi-Truths), a ConvNeXt baseline trained on DRCT taken from (Chen et al., 2024a), and an EfficientNet-B4 baseline trained on FaceForensics++ from (Yan et al., 2023). For semi-supervised detection, we include the CLIP-Based Synthetic Image Detector (Cozzolino et al., 2024), which applies a linear probe over CLIP embeddings with minimal training data. For zero-shot detection, we use InternVL (Chen et al., 2024c) directly without finetuning. [^4] We use the InternVL3-38B; Code for all scripts and baselines https://github.com/vicliv/OpenFake Finally, we also test the handcrafted detector of Corvi et al. (2023) and a hybrid fusion baseline that averages predictions from CLIP and Corvi2023. Together, these baselines cover a diverse range of architectures, training regimes, and prior benchmarks, allowing us to compare detectors trained on OpenFake against both legacy and contemporary approaches.
Results on OpenFake.
As shown in Table˜3, the SwinV2 trained on OpenFake is near-perfect, confirming that modern classifiers are highly reliable when trained on the evaluation distribution. The SwinV2 trained on GenImage is the next strongest and handles many shared open-source families, but degrades on newer proprietary generators (Grok 2, Midjourney 6, Flux-1.1 pro, DALL·E 3), reflecting a distribution gap in the fake images. The SwinV2 trained on Semi-Truths achieves high TPRs yet misclassifies many real images, which is consistent with its training data focused on edits rather than full generation. Legacy or narrow baselines (ConvNeXt/DRCT, EffNet-B4/FF++) underperform, CLIP-D-10k+ is middling, and Corvi2023 largely predicts “real,” while zero-shot InternVL is better than older baselines but still trails supervised models. Overall, robust performance requires training on the correct, broad, and up-to-date image distribution.
Metric / Model SwinV2 (Ours) SwinV2 (GenImage) SwinV2 (Semi-Truths) ConvNeXt (DRCT) EffNet-B4 (FF++) CLIP- D-10k+ Corvi 2023 InternVL-3 (zero-shot)
Training Dataset OpenFake GenImage (Zhu et al., 2023) Semi-Truths (Pal et al., 2024) DRCT (Chen et al., 2024a) FF++ (Rössler et al., 2019) Custom (Cozzolino et al., 2024) DMD (Corvi et al., 2023)
Real Images TNR 0.995 0.955 0.689 0.777 0.516 0.703 0.998 0.431
Stable Diffusion 1.5 1.000 0.936 1.000 0.447 0.529 0.579 0.000 0.849
Stable Diffusion 2.1 1.000 0.998 0.999 0.482 0.453 0.717 0.011 0.900
Stable Diffusion XL 1.000 0.956 1.000 0.426 0.507 0.438 0.001 0.814
Stable Diffusion 3.5 1.000 0.982 1.000 0.324 0.466 0.406 0.000 0.796
Flux 1.0 Dev 1.000 0.967 0.999 0.290 0.450 0.401 0.005 0.748
Flux-1.1-Pro 1.000 0.315 0.975 0.319 0.467 0.596 0.000 0.722
Flux-1.0-Schnell 0.999 1.000 0.998 0.289 0.476 0.503 0.000 0.803
Midjourney 6 1.000 0.090 0.949 0.166 0.486 0.100 0.000 0.884
Midjourney 7 0.994 0.952 0.997 0.264 0.484 0.404 0.001 0.961
DALL·E 3 0.995 0.238 0.927 0.461 0.543 0.394 0.000 0.983
GPT Image 1 0.998 0.772 0.983 0.402 0.442 0.384 0.005 0.932
Ideogram 3.0 1.000 0.993 1.000 0.254 0.481 0.414 0.001 0.844
Imagen 3.0 0.999 0.962 0.998 0.237 0.461 0.286 0.005 0.784
Imagen 4.0 0.996 0.948 0.996 0.228 0.459 0.359 0.003 0.796
Grok 2 1.000 0.142 0.963 0.383 0.463 0.303 0.000 0.805
HiDream-I1 Full 1.000 0.976 0.993 0.332 0.440 0.485 0.000 0.789
Chroma 0.992 0.980 0.995 0.451 0.435 0.298 0.003 0.726
Ideogram 2.0 0.993 0.997 1.000 0.234 0.482 0.777 0.000 0.865
Lumina 1.000 1.000 1.000 0.494 0.355 0.720 0.028 0.983
Frames 0.968 0.816 1.000 0.368 0.392 0.920 0.000 0.912
Halfmoon 0.995 0.953 1.000 0.263 0.353 0.632 0.000 0.832
Recraft v2 0.972 0.699 1.000 0.379 0.443 0.248 0.004 0.929
Recraft v3 0.701 0.288 0.997 0.364 0.497 0.430 0.002 0.912
Average TPR 0.988 0.823 0.992 0.354 0.475 0.443 0.003 0.827
Overall F1 0.992 0.881 0.861 0.449 0.485 0.509 0.005 0.697
Overall ROC AUC 1.000 0.926 0.960 0.616 0.493 0.600 0.487 0.629
Overall PR AUC 1.000 0.949 0.952 0.613 0.493 0.600 0.488 0.586
Table 3: Performance comparison on OpenFake across detectors trained on different datasets. Finetuned (FT) and LoRA variants are grouped under their respective base generators. Generators shown in blue are out-of-distribution for all detectors. SwinV2 trained on OpenFake consistently outperforms others on unseen generators, while most alternative detectors exhibit high false positive rates (misclassification of real images).
Test set Metric Train OpenFake Train GenImage Train Semi-Truths
GenImage TPR 0.771 1.000 0.965
TNR 0.928 1.000 0.261
Accuracy 0.849 1.000 0.613
F1 Score 0.836 1.000 0.714
Semi-Truths TPR 0.909 0.830 1.000
TNR 0.962 1.000 1.000
Accuracy 0.920 0.865 1.000
F1 Score 0.947 0.907 1.000
Table 4: Cross-benchmark generalization of SwinV2 detectors. Italicised numbers indicate in-domain evaluations, where the model is tested on the same dataset it was trained on. TPR = true-positive rate (synthetic images), TNR = true-negative rate (real images). All values are shown to three decimal places.
Importance of dataset coverage.
These results highlight two key takeaways. First, a well-trained supervised classifier like SwinV2 can detect even high-fidelity synthetic images if it has seen samples from that generator. Second, existing detection models trained on older benchmarks fail to generalize to unseen, state-of-the-art generators. This highlights the necessity of updated and evolving detection datasets, such as OpenFake, to match the capabilities of modern generative models.
Robustness to compression artifacts.
Real images (sourced from LAION-400M in our dataset) are typically compressed and carry authentic JPEG artifacts and blur. In contrast, synthetic images are high-resolution and minimally compressed. This mismatch creates an obvious signal: detectors may rely on compression differences instead of true semantic features, thus failing on compressed fakes. To assess this vulnerability, we implemented a data augmentation pipeline to degrade synthetic images during training, mimicking the distribution of real images, as done in previous work (Corvi et al., 2023; Wang et al., 2021; Cheng et al., 2024). This includes random resizing, Gaussian blur, JPEG compression, and Gaussian noise. We then evaluated the SwinV2 model trained with these augmentations on a fully compressed test set and got an overall F1 score of 0.992. This demonstrates that the model remains highly accurate even when the compression signal is neutralized.
Transferability to unseen models.
Using the out-of-distribution generators in Table˜3 (blue rows), which were collected from public web sources rather than generated by us, the SwinV2 trained on OpenFake shows the strongest transfer while keeping a high true-negative rate on real images. The SwinV2 trained on GenImage is competitive on several open-source families but lags on newer proprietary models, and the Semi-Truths model attains high TPRs yet mislabels many real images, so its apparent OOD gains are not reliable. Cross-benchmark tests in Table˜4 reinforce this: when evaluated on GenImage, the OpenFake model substantially outperforms the Semi-Truths model (Accuracy 0.849 vs. 0.613; F1 0.836 vs. 0.714), and when evaluated on Semi-Truths it exceeds the GenImage model (Accuracy 0.920 vs. 0.865; F1 0.947 vs. 0.907), despite being out-of-domain in both cases. Overall, the most transferable detector is the one trained on the broad, up-to-date distribution provided by OpenFake.
7 Conclusion
This paper introduces OpenFake, a politically grounded benchmark for evaluating deepfake detectors on high-fidelity, real-world-relevant synthetic imagery. We constructed the dataset by pairing three million politically themed real images filtered from LAION with over 963k synthetic counterparts generated using state-of-the-art open and proprietary models. To complement this static benchmark, we developed OpenFake Arena, a crowdsourced adversarial platform that incentivizes the continual creation of hard-to-detect synthetic images, enabling detectors to adapt to the evolving landscape of generative models. Through a human perception study, we showed that recent proprietary generators (e.g., Imagen 3, GPT Image 1) often fool users. We also demonstrated that detectors trained on older benchmarks fail to identify these modern fakes, while supervised models trained on OpenFake achieve strong performance, especially with compression-robust augmentations.
Our results highlight that training on the right, broad distribution is crucial. The SwinV2 model trained on OpenFake not only excels in-distribution but also transfers best to unseen generators collected from the web, outperforming baselines trained on narrower datasets. Generalization is not perfect, however; performance still varies across some proprietary or lower-quality sources, indicating that coverage and recency of training data matter.
Encouragingly, current detection methods can handle high-quality deepfakes when trained on appropriately curated and up-to-date data. OpenFake, with its real-world relevance, synthetic diversity, and extensibility through OpenFake Arena, offers a scalable path forward. By combining high-fidelity benchmarking with community-driven adversarial submissions, our framework aims to narrow the gap between generation and detection, equipping researchers and practitioners with tools to confront emerging misinformation threats in real time.
Acknowledgments and Disclosure of Funding
This work was partially funded by the CIFAR AI Chairs Program. Additional financial support was provided by the Centre for the Study of Democratic Citizenship (CSDC), IVADO, and the Canada First Research Excellence Fund. We also gratefully acknowledge Mila for both financial assistance and access to computational resources. Finally, we thank Fuxiao Gao and Jie Zang for their help with data collection, as well as all participants who contributed to the human perception survey.
Appendix A OpenFake composition and licensing
Table˜5 summarizes the dataset at the generator level, listing each base model and its LoRA or finetuned variants with release month, exact image counts, and license category. In total there are 963,342 synthetic images drawn from Stable Diffusion (1.5/2.1/XL/3.5), Flux (1.0 dev, 1.1 Pro, Schnell), Midjourney (6/7), DALL·E 3, Imagen (3/4), GPT Image 1, Grok 2, Ideogram 3.0, HiDream, Chroma, and Recraft v3. The real corpus contains 3M filtered LAION-400M images. Some proprietary and out-of-distribution generators appear with smaller totals because they were sourced from external collections rather than produced end-to-end.
Licensing and access are made explicit to support downstream compliance. We label sources as Community, Non-commercial, or Non-compete and include these labels in the release metadata. All manifests are hosted on the HuggingFace Hub in streaming-friendly formats (Parquet and CSV) with per-item metadata such as model family, variant, release month, and prompt text.
Source Release (YYYY–MM) # Images Licence
Real (LAION-400M, filtered) 2021–08 3M
Stable Diffusion 1.5 2022–08 76,510 Community
   Stable Diffusion 1.5 (base) 2022–08 20,000 Community
   Dreamshaper (FT) 2023–07 36,510 Community
   EpicDream (FT) 2023–08 20,000 Community
Stable Diffusion 2.1 2022–12 135,487 Community
Stable Diffusion XL 2023–07 186,666 Community
   Stable Diffusion XL (base) 2023–07 40,000 Community
   Epic Realism (FT) 2025–06 59,770 Community
   Touch of realism (LoRA) 2025–06 32,828 Community
   RealVisXL-v5 (FT) 2024–09 29,300 Community
   Juggernaut (FT) 2025–05 24,768 Community
Flux 1.0 dev 2024–08 144,788 Non-commercial
   Flux 1.0 dev (base) 2024–08 106,796 Non-commercial
   Mystic (FT) 2024–10 15,608 Non-commercial
   MVC5000 (LoRA) 2025–07 16,244 Non-commercial
   Amateur Snapshot Photos (LoRA) 2025–06 4,140 Non-commercial
   Realism (LoRA) 2024–08 2,000 Non-commercial
DALL·E 3 2023–10 33,336 Non-compete
Midjourney 6 2023–12 50,000 Non-compete
Imagen 3.0 2024–08 4,032 Non-compete
Flux-1.0-Schnell 2024–08 36,084 Non-commercial
Flux-1.1-Pro 2024–10 29,923 Non-commercial
Recraft v3 2024–10 1,000 Community
Stable Diffusion 3.5 2024–10 139,114 Non-compete
Grok 2 2024–12 9,803 Non-compete
Ideogram 3.0 2025–03 28,495 Non-compete
Midjourney 7 2025–04 3586 Non-compete
GPT Image 1 2025–04 41,315 Non-compete
HiDream-I1 Full 2025–04 27,904 Community
Imagen 4.0 2025–05 10,721 Non-compete
Chroma 2025–08 4,532 Community
Total synthetic 963,342
Table 5: OpenFake statistics. Image counts are exact. While we release the entire 3M real images and prompts, only a balanced subset is fully uploaded to the HuggingFace Hub to match the number of fake images. The remainder can be downloaded via URLs provided in CSV files on the Hub. LoRA variants (“LoRA”) and full finetunes (“FT”) are listed on separate, smaller rows directly below their base models.
Appendix B Filtering and Captioning of LAION Images
To curate a relevant subset of real images from LAION-400M, we implemented a two-stage filtering and captioning pipeline using the vision-language model Qwen2.5-VL. This approach allowed us to filter politically salient and emotionally impactful content while preserving real-world visual characteristics (e.g., compression artifacts) crucial for training robust deepfake detectors.
Filtering prompt.
The first step used a vision-language reasoning prompt to assess whether each image depicted (i) real human faces, and/or (ii) politically or emotionally significant events. Many original LAION captions are noisy or incomplete, so the model was asked to jointly analyze both image and caption. The prompt was:
Analyze the provided image and its caption: “{caption}”. Provide detailed reasoning on the following two points:
1. Does the image contain any real human face(s)? Exclude animations, cartoons, figurines, statues, drawings, paintings, or video games.
2. Does the image contain content related to political events, catastrophes, news events, or anything likely to have high emotional impact or polarization? Exclude animations, cartoons, drawings, paintings, or video games.
Conclude clearly with either “Humans: yes” or “Humans: no”, and “Catastrophes: yes” or “Catastrophes: no”.
Only images with at least one “yes” label (human or catastrophe) were retained. This strategy allowed us to target both portrait-based and event-based misinformation vectors while filtering out non-photographic and low-impact content.
Captioning prompt.
For the selected images, we generated improved prompts to guide synthetic image generation. These prompts describe the image in a style suitable for text-to-image models, incorporating visual format and subject matter. The Qwen2.5-VL prompt used was:
Given the image and its caption: “{caption}”, generate a concise prompt in a single sentence that describes the image and its format (e.g., photograph, poster, screenshot), including any people present. Do not mention the caption directly.
These refined prompts were used for synthetic image generation and are also included in the public release to support downstream research and reproducibility.
Appendix C Ethics, Privacy & Limitations
While our dataset aims to support robust deepfake detection, it inherits limitations from its sources. The real image corpus, derived from the LAION crawl (2014–2021), skews toward Western-centric and pre-pandemic imagery. Proprietary generative models also reflect aesthetic and cultural biases from their training data. These imbalances may affect the generalizability of detection models across diverse global contexts. We document these issues in the HuggingFace Data Card and encourage contributions from underrepresented regions via our Arena pipeline.
The paper includes details of both the human perception study and the Arena crowdsourcing platform. No compensation was offered, as participation was voluntary, and both systems were designed to ensure anonymity and avoid the collection of personal data.
Prompt extraction may introduce semantic noise, and the quality of adversarial data depends on user participation. Our dataset focuses on visual realism, but does not yet capture multimodal or context-based misinformation. Fairness across demographic groups and long-term robustness remain open challenges. We encourage downstream audits and broader evaluation to support responsible deployment.
Appendix D Training Details and Compute Resources
D.1 Compute Resources and Cost
All experiments were conducted on an internal compute cluster or local workstations with moderate storage and GPU availability. Below, we detail the computational resources and costs associated with dataset filtering, image generation, baseline evaluation, and dataset hosting.
Filtering and analysis.
The LAION filtering pipeline ran continuously for two weeks on 4 NVIDIA L40S GPUs (48 GB VRAM each). An additional 2 days of compute on the same setup was used for prompt selection and vision–language model evaluation, comparing multiple candidate models and prompt formats.
Synthetic image generation.
Images from Stable Diffusion v2.1 and Flux.1.0-dev were generated on 4 L40S GPUs over a span of 4 days per model. Other models generated images for 1 day. Each GPU was fully utilized to maximize throughput.
Model training and evaluation.
Training the SwinV2 baseline classifier on the OpenFake dataset required approximately 12 hours on a single NVIDIA L40S GPU. Inference for evaluation purposes was negligible in comparison.
Baseline inference.
For baseline evaluation:
  • InternVL inference over the full test set was performed over 10 hours on a single RTX8000 GPU (48 GB VRAM).
  • CLIP and the Corvi2023 baselines were evaluated in approximately 6 hours on the same RTX8000 GPU.
Proprietary model generation.
Images generated via proprietary APIs incurred a per-image cost of approximately \$0.04 (USD), varying slightly by model and resolution. No GPU compute was required on our end; generation was offloaded entirely to the remote API services.
Storage and hosting.
Dataset preprocessing, metadata formatting, and uploads to Hugging Face required only CPU cores but substantial storage capacity. The working set size during dataset preparation exceeded 1TB.
Total estimated GPU compute: $\sim$4 GPU-months across L40S and RTX8000 class cards. All compute was performed on institutional resources without incurring cloud costs.
D.2 SwinV2 Fine-Tuning Hyperparameters
For our main benchmark detector, we finetune microsoft/swinv2-small-patch4-window16-256 on the OpenFake dataset using the HuggingFace Trainer API. All experiments were conducted on a single L40S GPUs.
Model architecture.
We use the SwinV2-Small transformer backbone with the classifier head modified to predict two classes: real vs. fake. The model is initialized from ImageNet-1k weights and fine-tuned end-to-end.
Input resolution.
Images are resized to $256\times 256$ using the default SwinV2 image processor.
Training configuration.
  • Optimizer: AdamW
  • Learning rate: 5e-5
  • Batch size: 32
  • Epochs: 5
  • Learning rate scheduler: Linear with warmup
Data augmentation.
During training we use two augmentation streams. A general geometric/photometric stream is applied to both real and synthetic images, including random resized crops, color jitter, small rotations, occasional horizontal flips, and mild Gaussian blur. To neutralize compression shortcuts, a light degradation stream is applied to synthetic images only, including resolution downscaling, blur adjustment, low-level Gaussian noise, and JPEG compression with randomized quality. Transforms are sampled stochastically, and the synthetic-only degradations are calibrated to match statistics of LAION-derived real images. For compressed test-set evaluation, synthetic images are post-processed with the same degradation function to simulate internet-style artifacts; we report accuracy, precision, recall, F1, and ROC AUC.
D.3 Generation Parameters for Open-Source Models
We document here the generation settings used to produce synthetic images from open-source models within the OpenFake dataset. This ensures reproducibility and clarity on the diversity of generated outputs.
We used stabilityai/stable-diffusion-3.5-large to generate synthetic images and black-forest-labs/Flux.1.0-dev using the same bank of prompts. Both models were run in bfloat16 precision using their official pipelines—StableDiffusion3Pipeline and FluxPipeline, respectively—and deployed across multiple GPUs with prompt sharding and batched inference for scalability. We used the official HugginFace weights for the other models via the Diffusers Python library.
For all models, the following generation settings were generally applied (there could be slight modifications based on the recommended parameters for each model):
  • Resolution: Randomly sampled from a predefined set of social-media-style sizes: [(1024, 1024), (1024, 512), (512, 1024), (1024, 768), (768, 1024), (1152, 768), (768, 1152)]
  • Guidance scale: Uniformly sampled between 1.5 and 7
  • Inference steps: [10, 40]
  • Scheduler: Default
These configurations were chosen to maximize diversity and photorealism, while reflecting the resolution and stylistic variability typical of online content.
Appendix E OpenFake Arena
We host the Arena as a Gradio app on Hugging Face Spaces, leveraging their compute resources. A pretrained CLIP model acts as a prompt-matching gate to ensure image relevance, and successful submissions that fool the detector are stored in a connected Hugging Face dataset. The detector is a SwinV2 model trained on the OpenFake dataset and periodically updated to reflect new data. We also log metadata such as the generative model used and the user ID to support leaderboard tracking. Prompts are designed to be specific and difficult to spoof, and additional safeguards are in place to prevent misuse. Upon acceptance, we plan to promote the Arena through social media and at the conference to encourage broader participation. Figures 5, 7, and 6 show the Arena interface and leaderboard, along with examples of successful and failed submissions.
Figure 5: OpenFake Arena interface. Users are presented with a prompt and asked to generate an image that can fool the detector.
Figure 6: Example of a successful submission. The image aligns with the prompt "The cover of "Erased" by Jennifer Rush features a dramatic image of a shirtless man with a stormy sky and lightning, set against a grid background, creating a suspenseful atmosphere". It is incorrectly classified as real by the detector.
Figure 7: Example of an unsuccessful submission. The image fails to fool the detector and is correctly classified as synthetic. The prompt used was "A photograph captures Dianne Reeves performing on stage in the East Room of the White House during the National Governors Association Dinner on February 26, 2012, with an audience seated in the foreground."
Appendix F Synthetic image examples from OpenFake
[…]
Figure 8: Sample images from OpenFake generated by GPT Image 1.
[…]
Figure 9: Sample images from OpenFake generated by Ideogram 3.0.
[…]
Figure 10: Sample images from OpenFake generated by Flux-1.1-Pro.
[…]
Figure 11: Sample images from OpenFake generated by Flux.1-Dev.
[…]
Figure 12: Sample images from OpenFake generated by Stable Diffusion 3.5.